# Bridging the classical and quantum regimes in a dissipative Ising chain

### v1

@@ VI Summary
In summary, we propose an ensemble averaged cluster model with an interpolated coherent Hamiltonian to study the long-time dynamics of a dissipative Ising chain. Starting from the classical limit where the long-time dynamics exhibit either limit cycles or near chaotic behaviors, we show how the system sequentially undergoes frequency bifurcations and a steady-state phase transition, so that the persistent oscillations of the limit cycles change into steady-state approaching dynamics. Our model is relevant to dissipative Rydberg gases, wherein the interpolation parameter $\lambda$, which qualitatively measures the quantum correlations within the system, could be tuned through the temperature. For future studies, it would be desirable to apply a similar approach to dissipative Ising models in higher dimensions.

### v2

@@ VI Summary
In summary, we propose an ensemble averaged cluster model with an interpolated coherent Hamiltonian to study the long-time dynamics of a dissipative Ising chain. Starting from the classical limit where the long-time dynamics exhibit either limit cycles or near chaotic behaviors, we show how the system sequentially undergoes frequency bifurcations and a steady-state phase transition, so that the persistent oscillations of the limit cycles change into steady-state approaching dynamics. Our model is relevant to dissipative Rydberg gases, wherein the interpolation parameter $\lambda$, which qualitatively measures the quantum correlations within the system. For future studies, it would be desirable to apply a similar approach to dissipative Ising models in higher dimensions.
@@ Acknowledgements.
Data Availability
The data that support the findings of this study are openly available [60].